OpenStudy (anonymous):

find fraction notation and simplify 12.5% SHOW WORK PLEASE

OpenStudy (michele_laino):

\[=\frac{ 12.5 }{ 100 }=\frac{ 25/2 }{ 100 }=\frac{ 25 }{ 200 }=\frac{ 1 }{ 8 }\]

OpenStudy (michele_laino):

@cucu

OpenStudy (anonymous):

what about 33 1/3%

OpenStudy (lizz123):

http://fractionsimplifier.com/mixed-number-simplifier/ It shows what it does

OpenStudy (lizz123):

It shows the works under results if you scroll down

OpenStudy (lizz123):

@cucu